# On the mechanism of black hole energy reduction in the Blandford-Znajek process

### v1

@@ Abstract
The Blandford-Znajek (BZ) process is electromagnetic energy release from rotating black holes (BHs) along magnetic field lines threading them and widely believed to drive relativistic jets. This process is successfully demonstrated in general relativistic magnetohydrodynamic (MHD) simulations with the coordinate system regular on the event horizon, by which one can estimate the outward Poynting flux, although the direct energy release through the horizon shown in the simulations does not provide an intuitive picture. We revisit the mechanism of BH energy reduction by utilizing the coordinate system singular on the horizon, in which the falling membrane of past accreted matter should exist above the horizon. We find that the Poynting flux is produced at the boundary between the falling membrane and the magnetically-dominated inflow, and the front of the inflow creates the negative electromagnetic energy, which reduces the rotational energy of spacetime. We also clarify that the poloidal electric current does not form a closed circuit within the magnetically-dominated flow. Previous interpretations of the BZ process and possibilities of the ideal MHD violation and BH charging are also discussed.
@@ 1 Introduction
Relativistic jets are collimated outflows with velocities close to the light speed, which are observed in active galactic nuclei, gamma-ray bursts, galactic microquasars, and so on. Their main driving mechanism is widely believed to be the Blandford-Znajek (BZ) process blandford77 , i.e., the electromagnetic extraction of black hole (BH) rotational energy. General relativistic magnetohydrodynamic (GRMHD) simulations have been developed since 2000s komissarov01 ; koide02 ; komissarov04b , and the simulations of radiatively inefficient accretion flows (RIAFs) have shown that BH magnetospheres can form, where outward Poynting flux is maintained along the ordered magnetic field lines threading the event horizon via the BZ process and fluid is electromagnetically accelerated to the relativistic speed (e.g. mckinney04, ; barkov08, ; tchekhovskoy11, ; takahashi16, ; nakamura18, ; porth19, ; liska20, ; narayan22, ; ripperda22, ). This picture is supported by very long baseline interferometric observations of the jet of M87, which is one of the closest radio galaxies (e.g. nakamura18, ; hada11, ; eht19a, ; eht19b, ; eht21, ; eht24, ).
@@
The BZ process was theoretically discovered by analogy with the unipolar induction process of rotating neutron stars with dipole magnetic field blandford77 ; goldreich69 ; wald74 . In the latter case, the neutron star is matter-dominated, where the poloidal electric current is driven towards the higher electric potential by the rotation of matter, producing the Poynting flux and reducing the rotational energy of matter beskin10 (see also shibata21, ). In contrast, there is no matter-dominated region in the BZ process. The Poynting flux appears to directly reduce the BH rotational energy without effective exchange between matter and electromagnetic energies. This is clearly demonstrated in the GRMHD simulations with coordinate system regular on the event horizon, Kerr-Schild (KS) coordinates. [^1] The GRMHD simulations fix the spacetime metric. The decrease of BH spin is actually shown by solving Einstein equation with an analytic perturbative method kimura21 and with numerical relativity simulations shibata24 .
@@
However, the BH energy reduction by the direct energy release from the event horizon does not give us an intuitive picture. The outward Poynting flux on the horizon is possible because this flux vector is space-like in the ergosphere, where no physical observer measures this flux komissarov09 ; lasota14 . This flux is mathematically conserved along each magnetic field line (under the force-free approximation), and measurable by an observer at the far zone. Some theorists thus tried to find an alternative picture that invokes ‘inflow of negative energy objects’ in a similar manner to the mechanical Penrose process penrose69 , but all of the proposed definitions of the negative energy and/or its inflow velocity are ambiguous komissarov09 ; koide14 ; kinoshita18 (see Section 5).
@@
In this paper, we provide a clear, intuitive picture for the mechanism of BH energy reduction in the BZ process by using the cooridate system singular on the event horizon, Boyer-Lindquist (BL) coordinates. A novel point in our study is that we treat a falling membrane of past accreted matter that should exist above the horizon. Then we find that the Poynting flux is produced outside the event horizon, associated with production of negative electromagnetic energy. We also clarify that the poloidal electric current that maintains the Poynting flux cannot form a closed circuit within the magnetosphere. Such a clarification may be useful for gaining insights into detailed plasma physics close to the horizon.
@@
In Section 2, we briefly review the description of particles and electromagnetic fields in the Kerr spacetime and the force-free split-monopole solution of the BZ process. In Section 3, we show an analytic solution of steady, axisymmetric GRMHD inflow along a magnetic field line. Based on these knowledges, we provide our picture for the BH energy reduction and the electric current closure in Section 4. We discuss the previous interpretations of the BZ process and current driving force in Section 5.
@@ 2.1 The Kerr spacetime and particles near the horizon
The ergosphere is the region of $r<1+\sqrt{1-a^{2}\cos^{2}\theta}$, where the Killing vector $\xi^{\mu}=(1,0,0,0)$ is space-like, $\xi^{2}=g_{tt}=-\alpha^{2}+\gamma_{\varphi\varphi}\Omega^{2}>0$. Thus, the world line of constant $(\varphi,r,\theta)$ is space-like in the ergosphere, and time-like particle motions are restricted to $d\varphi>0$ (‘frame-dragging effect’). The particle energy per mass $-u_{\mu}\xi^{\mu}=-u_{t}$ can be negative in the ergosphere. The reduction of BH rotational energy via infalling of a particle with $-u_{t}<0$ is called mechanical Penrose process penrose69 .
@@
by integration of geodesic equation carter68 ; punsly08 . This corresponds to $v^{\hat{r}}\to-1$ in the FIDO frame, which results from the divergently strong gravitational force (proportinal to $\partial_{r}\ln\alpha$ thorne86 ; levinson18 ). The energy per mass of the particle diverges as $-u_{\hat{t}}=-u_{\mu}e^{\mu}_{(t)}=-(u_{t}+\Omega u_{\varphi})/\alpha\propto\alpha^{-1}$, where we note that $u_{t}=u_{\mu}\xi^{\mu}$ and $u_{\varphi}=u_{\mu}\chi^{\mu}$ are constant along the geodesic.
@@ 2.2 The 3+1 electrodynamics
where $\nabla\cdot\mathbf{C}$ and ×$\nabla\times\mathbf{C}$ denote $(1/\sqrt{\gamma})\partial_{i}(\sqrt{\gamma}C^{i})$ and $(1/\sqrt{\gamma})\epsilon^{ijk}\partial_{j}C_{k}$, respectively, and $\epsilon^{ijk}$ is the Levi-Civita pseudo-tensor. $D^{\mu}$ and $B^{\mu}$ are the electric and magnetic fields as measured by the FIDOs, while $E^{\mu}$ and $H^{\mu}$ are those in the coordinate basis. These sets of electomagnetic fields have constitutive relations,
@@
where $\mathbf{m}=\partial_{\varphi}$. The charge density and the electric current in Eq. (8) are defined by $\rho_{e}=-I^{\mu}n_{\mu}$ and $J^{\mu}=(\xi^{\mu}I^{\nu}-\xi^{\nu}I^{\mu})n_{\nu}$.
@@
The angular momentum density and the angular momentum flux are written by $\ell~{}(=\alpha T^{t}_{~{}\varphi})=(1/4\pi)(\mathbf{D}\times\mathbf{B})\cdot\mathbf{m}$ and $L^{i}~{}(=\alpha T^{i}_{~{}\varphi})=-(\mathbf{E}\cdot\mathbf{m})\mathbf{D}-(\mathbf{H}\cdot\mathbf{m})\mathbf{B}+(1/2)(\mathbf{E}\cdot\mathbf{D}+\mathbf{B}\cdot\mathbf{H})\mathbf{m}$, respectively.
@@
2.3 Force-free split-monopole BZ solution
Blandford and Znajek blandford77 consider steady, axisymmetric magnetospheres of Kerr BHs, in which the event horizon is threaded by the ordered magnetic field produced by the electric current of plasma, and the energy density of electromagnetic field dominates that of the plasma particles. The latter condition (‘force-free condition’) is equivalent to no energy exchange between the electromagnetic field and plasma particles ($F_{\mu\nu}I^{\nu}=0$), that is,
$\rho_{e}\mathbf{E}+\mathbf{J}\times\mathbf{B}=0.$ (15)
The force-free condition (Eq. 15) means $\mathbf{E}\cdot\mathbf{B}=0$. For the steady, axisymmetric electric field, one has $E_{\varphi}=0$ from ×$\nabla\times\mathbf{E}=0$ (in Eq. 7), so that the electric field can be written by
@@ 2.3 Force-free split-monopole BZ solution
Then the energy and angular momentum equations (Eqs. 10 and 11) mean that $H_{\varphi}$ is constant along each magnetic field line.
In the slow rotation limit $a\ll 1$ with split-monopole poloidal magnetic field at the zeroth order ($B^{r}\propto 1/\sqrt{\gamma}$ and $B^{\theta}=B^{\varphi}=0$), Eq. (15) can be analytically solved to obtain (blandford77, )
$\Omega_{\rm F}=\frac{1}{2}\Omega_{\rm H}+\mathcal{O}(a^{3}).$ (18)
For deriving this solution, the two conditions are utilized; the regularity condition at the event horizon (Eq. 14),
$H_{\varphi}=(\Omega_{\rm F}-\Omega_{\rm H})\sqrt{\gamma}B^{r}\sin\theta,$ (19)
and the condition at the far zone michel73 , $\hat{B}_{\varphi}=\hat{E}_{\theta}$, which is equivalent to
$H_{\varphi}=-\Omega_{\rm F}\sqrt{\gamma}B^{r}\sin\theta.$ (20)
The zeroth order solution $\Omega_{\rm F}\approx\Omega_{\rm H}/2$ can be derived by matching the two conditions along each field line with the constancy of $H_{\varphi},\Omega_{\rm F},\sqrt{\gamma}B^{r},$ and $\theta$ (komissarov04a, ).
@@ 3 Steady axisymmetric MHD flows
[…]
Under the ideal MHD condition, $F_{\mu\nu}u^{\nu}=0$, one has $\mathbf{E}\cdot\mathbf{B}=0$ similarly to the force-free condition, so that $\Omega_{\rm F}$ is defined by Eq. (16) for steady, axisymmetric flows. One also has $\mathbf{E}=-\mathbf{v}\times\mathbf{B}$, which is reduced to
@@
where $\kappa$ is a scalar function. In addition to $\Omega_{\rm F}$, one has three quantities conserved along each magnetic field line,
@@
where $\rho$ is the fluid-frame mass density bekenstein78 . Given the poloidal field profile, $B^{r}$ and $B^{\theta}$, with the conserved quantities, $\Omega_{\rm F},\mathcal{E},$ and $\mathcal{L}$ for each field line, the above equations can be solved. ($\eta$ does not affect the structure of flows.) The flow solution consists of inflow ($\kappa<0,\eta<0,\mathcal{E}<0,\mathcal{L}<0$) and outflow ($\kappa>0,\eta>0,\mathcal{E}>0,\mathcal{L}>0$) which are separated at a surface where the gravity and Lorentz force balances. $\Omega_{\rm F},\mathcal{E},$ and $\mathcal{L}$ are determined by the regularity condition at the inner magnetosonic point in the inflow, that at the outer magnetosonic surface in the outflow, and the location of the separation surface takahashi90 ; pu20 . To find the poloidal field profile, one must solve the transverse force balance beskin00 ; huang20 ; ogihara21 .
Figure 1: Solution of steady, axisymmetric MHD inflow towards the event horizon of a Kerr BH along a magnetic field line. The parameters are $a=0.8$, $\Omega_{\rm F}=\Omega_{\rm H}/2$, $\mathcal{E}=-10$, and $\mathcal{L}\Omega_{\rm F}=-10.55$. The top left panel shows $M^{2}=4\pi\rho\alpha u_{\rm p}^{2}/B_{\rm p}^{2}$ with $-f$ and $-f+\beta^{2}$ as functions of $r-r_{\rm H}$, while the top right panel $H_{\varphi}/4\pi\eta$, $-u_{t}$, and $u_{\varphi}$. For the bottom panels, the poloidal magnetic field is assumed to be in the split-monopole shape, i.e., $B^{\theta}=v^{\theta}=0$. The bottom left panel shows $-v^{r}$ and $v^{\varphi}$ with $\alpha/\sqrt{\gamma_{\varphi\varphi}}$, $\Omega$, and $\Omega_{\rm F}$ as functions of $r-r_{\rm H}$, while the bottom right panel $B_{\hat{r}}/4\pi\eta$, $B_{\hat{\varphi}}/4\pi\eta$, $-D_{\hat{\theta}}/4\pi\eta$, and $e/20\pi\eta^{2}$.
Here we present an analytic solution of the inflow along a field line by using the method of Takahashi and Tomimatsu takahashi08 (see also pu20, ). Combining Eqs. (21), (22), (23), and (24) with $u_{\mu}u^{\mu}=-1$ leads to the Berunoulli equation,
@@
where $H_{\varphi}^{2}=H^{\varphi}H_{\varphi}$ and $B_{\rm p}^{2}=B^{r}B_{r}+B^{\theta}B_{\theta}$. (See Appendix A for derivation of Eq. (25).) The method of takahashi08 assumes that the function $\beta^{2}$ is regular at the inner magnetosonic point, so that the regularity condition of the trans-magnetosonic flow solution is automatically satisfied. A convenient way is to set the form of $x\equiv D_{\rm p}^{2}/B_{\varphi}^{2}=\gamma_{\varphi\varphi}(\Omega_{\rm F}-\Omega)^{2}/\beta^{2}$ as
$x=\left(1+C\frac{\Delta}{\Sigma}\right)\left(\frac{\Omega-\Omega_{\rm F}}{\Omega_{\rm H}-\Omega_{\rm F}}\right)^{2},$ (28)
@@
where we may take the numerical constant $C=-1$. Even for the other values of $C$, we find that the solution hardly changes. This form of $x$ was found to be consistent with that seen in a numerical solution that takes account of the transverse force balance ogihara21 .
@@
Figure 1 shows an inflow solution. The parameters for this solution are $\Omega_{\rm F}=\Omega_{\rm H}/2$, $\mathcal{E}=-10$, and $\mathcal{L}\Omega_{\rm F}=-10.55$. The BH spin parameter is set to be $a=0.8$, for which the horizon radius and BH angular velocity are $r_{\rm H}\approx 1.6$ and $\Omega_{\rm H}\approx 0.25$, respectively. The upper left panel of Figure 1 shows Alfvén Mach number $M^{2}=4\pi\rho\alpha u_{\rm p}^{2}/B_{\rm p}^{2}$ as a function of $r-r_{\rm H}$. The inflow starts at the separation point, $r\simeq r_{\rm H}+1.3$, passes the inner Alfvén point where $M^{2}=-f$ and the inner magnetosonic point where $M^{2}=-f+\beta^{2}$, and continues towards the horizon.

### v2

@@ Abstract
The Blandford-Znajek (BZ) process is steady electromagnetic energy release from rotating black holes (BHs) along magnetic field lines threading them and widely believed to drive relativistic jets. This process is successfully demonstrated in general relativistic magnetohydrodynamic (MHD) simulations with the coordinate system regular on the event horizon, in which the outward Poynting flux on the horizon is considered to reduce BH energy. Meanwhile, alternative pictures for the BH energy reduction that invoke infall of negative energy objects were also discussed, although all of the proposed definitions of the negative energy and/or its infall velocity were ambiguous. We revisit the mechanism of BH energy reduction in the BZ process under the ideal MHD condition by utilizing the coordinate system singular on the horizon, in which the falling membrane of past accreted matter should exist above the horizon. We find that the Poynting flux is produced at the boundary between the falling membrane and the magnetically-dominated inflow, and the front of the inflow creates the negative electromagnetic energy, which reduces the rotational energy of spacetime. We also clarify that the poloidal electric current does not form a closed circuit within the magnetically-dominated flow. Previous interpretations of the BZ process and possibilities of violation of ideal MHD condition and BH charging are also discussed.
@@ 1 Introduction
Relativistic jets are collimated outflows with velocities close to the light speed, which are observed in active galactic nuclei, gamma-ray bursts, galactic microquasars, and so on. Their main driving mechanism is widely believed to be the Blandford-Znajek (BZ) process blandford77 , i.e., the steady electromagnetic extraction of black hole (BH) rotational energy. General relativistic magnetohydrodynamic (GRMHD) simulations have been developed since 2000s komissarov01 ; koide02 ; komissarov04b , and the simulations of radiatively inefficient accretion flows (RIAFs) have shown that BH magnetospheres can form, where outward Poynting flux is maintained along the ordered magnetic field lines threading the event horizon via the BZ process and fluid is electromagnetically accelerated to the relativistic speed (e.g. mckinney04, ; barkov08, ; tchekhovskoy11, ; takahashi16, ; nakamura18, ; porth19, ; liska20, ; narayan22, ; ripperda22, ). This picture is supported by very long baseline interferometric (VLBI) observations of the jet of M87, which is one of the closest radio galaxies (e.g. nakamura18, ; hada11, ; eht19a, ; eht19b, ; eht21, ; eht24, ).
@@
The BZ process was theoretically discovered by analogy with the unipolar induction process of rotating neutron stars with dipole magnetic field blandford77 ; goldreich69 ; wald74 . The neutron star is matter-dominated, where the poloidal electric current is driven towards the higher electric potential by the rotation of matter, producing the Poynting flux and reducing the rotational energy of matter beskin10 (see also shibata21, ). In contrast, there is no matter-dominated region in the BZ process. The Poynting flux appears to directly reduce the BH rotational energy without effective exchange between matter and electromagnetic energies. This is clearly demonstrated in the GRMHD simulations with coordinate system regular on the event horizon, Kerr-Schild (KS) coordinates. [^1] The GRMHD simulations fix the spacetime metric. The decrease of BH spin is actually shown by solving Einstein equation with an analytical perturbative method kimura21 and with numerical relativity simulations shibata24 .
@@
The outward Poynting flux on the horizon in the GRMHD simulations is possible because this flux vector is space-like in the ergosphere, and no physical observer measures this flux komissarov09 ; lasota14 . The outward Poynting flux is measurable by an observer at the far zone. Some theorists tried to find an alternative picture that invokes infall of negative energy objects in a similar manner to the mechanical Penrose process penrose69 , but all of the proposed definitions of the negative energy and/or its infall velocity were ambiguous takahashi90 ; komissarov09 ; koide14 (see Section 5.1.3).
@@
In this paper, we provide a clear picture for the mechanism of BH energy reduction with negative electromagnetic energy in the BZ process by using the cooridate system singular on the event horizon, Boyer-Lindquist (BL) coordinates. For the basis of this picture, we show analytical GRMHD solutions along a magnetic field line near the horizon. A novel point in our picture is that we treat a falling membrane of past accreted matter that should exist above the horizon. Then we find that the Poynting flux is produced outside the event horizon, associated with production of negative electromagnetic energy. We also clarify that the poloidal electric current that maintains the Poynting flux cannot form a closed circuit within the magnetosphere. Such a clarification may be useful for gaining insights into detailed plasma physics close to the horizon.
@@
In Section 2, we briefly review the description of particles and electromagnetic fields in the Kerr spacetime and the formulation of steady, axisymmetric Kerr BH magnetospheres. In Section 3, we show analytical solutions of steady, axisymmetric GRMHD inflow along a magnetic field line. Based on these knowledges, we provide our picture for the BH energy reduction and the electric current closure in Section 4. We discuss the previous interpretations of the BZ process and the driving force of poloidal current in Section 5.
@@ 2.1 The Kerr spacetime and particles near the horizon
The ergosphere is the region of $r<1+\sqrt{1-a^{2}\cos^{2}\theta}\equiv r_{\rm es}$, where the Killing vector $\xi^{\mu}=(1,0,0,0)$ is space-like, $\xi^{2}=g_{tt}=-\alpha^{2}+\gamma_{\varphi\varphi}\Omega^{2}>0$. Thus, the world line of constant $(\varphi,r,\theta)$ is space-like in the ergosphere, and time-like particle motions are restricted to $d\varphi>0$ (‘frame-dragging effect’). The particle energy per mass $-u_{\mu}\xi^{\mu}=-u_{t}$ can be negative in the ergosphere. The reduction of BH rotational energy via infalling of a particle with $-u_{t}<0$ is called mechanical Penrose process penrose69 .
@@
by integration of geodesic equation carter68 ; punsly08 . This corresponds to $v^{\hat{r}}\to-1$ in the FIDO frame, which results from the divergently strong gravitational force (proportional to $\partial_{r}\ln\alpha$ thorne86 ; levinson18 ). The energy per mass of the particle diverges as $-u_{\hat{t}}=-u_{\mu}e^{\mu}_{(t)}=-(u_{t}+\Omega u_{\varphi})/\alpha\propto\alpha^{-1}$, where we note that $u_{t}=u_{\mu}\xi^{\mu}$ and $u_{\varphi}=u_{\mu}\chi^{\mu}$ are constant along the geodesic.
@@ 2.2 The 3+1 electrodynamics
where $\nabla\cdot\mathbf{C}$ and ×$\nabla\times\mathbf{C}$ denote $(1/\sqrt{\gamma})\partial_{i}(\sqrt{\gamma}C^{i})$ and $(1/\sqrt{\gamma})\epsilon^{ijk}\partial_{j}C_{k}$, respectively, and $\epsilon^{ijk}$ is the Levi-Civita pseudo-tensor. $D^{\mu}$ and $B^{\mu}$ are the electric and magnetic fields as measured by the FIDOs, while $E^{\mu}$ and $H^{\mu}$ are those in the coordinate basis. These sets of electromagnetic fields have constitutive relations,
@@
where $\mathbf{m}=\partial_{\varphi}$ and $\mathbf{C}\times\mathbf{F}$ denotes $(1/\sqrt{\gamma})\epsilon^{ijk}C_{i}F_{k}$. The charge density and the electric current in Eq. (8) are defined by $\rho_{e}=-I^{\mu}n_{\mu}$ and $J^{\mu}=(\xi^{\mu}I^{\nu}-\xi^{\nu}I^{\mu})n_{\nu}$.
@@
The angular momentum density and the angular momentum flux are written by $\ell~{}(=\alpha T^{t}_{~{}\varphi})=(1/4\pi)(\mathbf{D}\times\mathbf{B})\cdot\mathbf{m}$ and $L^{i}~{}(=\alpha T^{i}_{~{}\varphi})=-(1/4\pi)(\mathbf{E}\cdot\mathbf{m})\mathbf{D}-(1/4\pi)(\mathbf{H}\cdot\mathbf{m})\mathbf{B}+(1/8\pi)(\mathbf{E}\cdot\mathbf{D}+\mathbf{B}\cdot\mathbf{H})\mathbf{m}$, respectively.
@@
2.3 Steady, axisymmetric magnetospheres and the BZ solution
In this paper we basically consider steady, axisymmetric magnetospheres of Kerr BHs, in which the electromagnetic energy density dominates the plasma particle energy density, and the event horizon is threaded by the ordered magnetic field produced by the electric current of plasma. We consider the BZ process as the steady production of outward Poynting flux along such ordered magnetic field lines. [^2] It can be generically proved that the BZ process operates, i.e., $\Omega_{\rm F}>0$ and $H_{\varphi}\neq 0$, when the magnetic field lines are threading the ergosphere and $\mathbf{D}\cdot\mathbf{B}=0$ (komissarov04a, ; toma14, ). Toma and Takahara (toma14, ) argued that the field lines need to cross the outer light surface, but only crossing of inner light surface is sufficient.
We assume that the particle number density is high enough to screen the electric field along the magnetic field lines,
$\mathbf{D}\cdot\mathbf{B}=0.$ (15)
This condition and Eq. (9) lead to $\mathbf{E}\cdot\mathbf{B}=0$. In the steady state, we have ×$\nabla\times\mathbf{E}=0$, which means that $\mathbf{E}$ is a potential field, and the axisymmetry leads to $E_{\varphi}=0$. Then one can write
@@ 2.3 Steady, axisymmetric magnetospheres and the BZ solution
The angular momentum equation (Eq. 11), $\nabla\cdot(-H_{\varphi}\mathbf{B}_{\rm p}/4\pi)=B^{i}\partial_{i}(-H_{\varphi}/4\pi)=-(\mathbf{J}_{\rm p}\times\mathbf{B}_{\rm p})\cdot\mathbf{m}$, implies that $H_{\varphi}={\rm const.}$ along a magnetic field line if the poloidal electric current does not cross it, $\mathbf{J}_{\rm p}\times\mathbf{B}_{\rm p}=0$.
Based on the above formulation, Blandford and Znajek blandford77 further assume the force-free condition, $\rho\mathbf{E}+\mathbf{J}\times\mathbf{B}=0$, to find perturbative solutions of electromagnetic field in the slow rotation limit $a\ll 1$. For deriving the solutions, the regularity condition at the event horizon (Eq. 14) and the condition at the far zone michel73 , $B_{\hat{\varphi}}=E_{\hat{\theta}}$, are utilized. They are written by $H_{\varphi}=(\Omega_{\rm F}-\Omega_{\rm H})\sqrt{\gamma}B^{r}\sin\theta$ and $H_{\varphi}=-\Omega_{\rm F}\sqrt{\gamma}B^{r}\sin\theta$, respectively, in the case of split-monopole magnetic field at the zeroth order ($B^{r}\propto 1/\sqrt{\gamma}$ and $B^{\theta}=B^{\varphi}=0$). The solution $\Omega_{\rm F}\approx\Omega_{\rm H}/2$ can be understood by matching the two conditions along each field line with the constancy of $H_{\varphi},\Omega_{\rm F},\sqrt{\gamma}B^{r},$ and $\theta$ (komissarov04a, ).
It was argued that the BZ solutions of electromagnetic field with the force-free condition have a causality problem, since they are determined by the conditions at the event horizon and at infinity punsly89 . The horizon must not yield any condition that actively affects its exterior, only passively absorbing particles and waves. This causality problem is avoided in the approach with the MHD condition takahashi90 ; levinson04 ; komissarov04a ; beskin10 (see Sections 3 and 4).
@@ 3 Steady axisymmetric MHD flows
In this section we investigate the detailed structure of steady, axisymmetric, cold flows in the BZ process under the ideal MHD condition, based on the formalism presented in Section 2.3. This is required for building a clear picture for the mechanism of BH energy reduction (Section 4).
The ideal MHD condition, $F_{\mu\nu}u^{\nu}=0$, is reduced to $\mathbf{E}=-\mathbf{v}\times\mathbf{B}$, so that $\mathbf{E}\cdot\mathbf{B}=0$, and $\Omega_{\rm F}$ is defined by Eq. (16) for steady, axisymmetric flows. Then one has
@@
where $\kappa$ is a scalar function. In addition to $\Omega_{\rm F}$, one has three quantities conserved along each magnetic field line from the conservation equations of mass, energy, and angular momentum bekenstein78 ,
@@
where $\rho$ is the fluid-frame mass density. Given the poloidal field profile, $B^{r}$ and $B^{\theta}$, with the conserved quantities, $\Omega_{\rm F},\mathcal{E},$ and $\mathcal{L}$ for each field line, the above equations can be solved for $u_{\mu}$ and $H_{\varphi}$. $\eta$ does not affect the structure of flows (see Appendix A). The flow solution consists of inflow ($\kappa<0,\eta<0,\mathcal{E}<0,\mathcal{L}<0$) and outflow ($\kappa>0,\eta>0,\mathcal{E}>0,\mathcal{L}>0$) which are separated at the point where the gravity and Lorentz force are balanced.
$\Omega_{\rm F},\mathcal{E},$ and $\mathcal{L}$ are determined by the regularity condition at the inner magnetosonic point in the inflow, that at the outer magnetosonic point in the outflow, and the location of the separation point, without the condition at the event horizon takahashi90 ; pu20 . For an inappropriate set of $\Omega_{\rm F},\mathcal{E},$ and $\mathcal{L}$, one cannot obtain the trans-magnetosonic inflow continuous towards the horizon through the inner magnetosonic point or the trans-magnetosonic outflow continuous towards infinity through the outer magnetosonic point. To find the poloidal field profile, one must solve the transverse force balance between the magnetic field lines beskin00 ; huang20 ; ogihara21 .
Here we focus on the inflow and present its analytical solutions along a field line in the BL coordinates by using the method of Takahashi and Tomimatsu takahashi08 (see also pu20, ). We do not solve the transverse force balance. Combining Eqs. (18), (19), (20), and (21) with $u_{\mu}u^{\mu}=-1$ leads to the Bernoulli equation,
@@
where $H_{\varphi}^{2}=H^{\varphi}H_{\varphi}$ and $B_{\rm p}^{2}=B^{r}B_{r}+B^{\theta}B_{\theta}$ (see Appendix A for derivation of Eq. 22). For given $\mathbf{B}_{\rm p}$, the appropriate set of $\Omega_{\rm F},\mathcal{E},$ and $\mathcal{L}$ must be searched for obtaining the trans-magnetosonic flow solution as explained above, whereas the method of takahashi08 assumes that the function of $\beta^{2}$ is regular, for which one can obtain the trans-magnetosonic solution and $\mathbf{B}_{\rm p}$ for a given set of $\Omega_{\rm F},\mathcal{E},$ and $\mathcal{L}$. In this method the regularity condition of the flow solution at the inner magnetosonic point is automatically satisfied just by the assumption on $\beta^{2}$. A convenient way is to set the form of $x\equiv D_{\rm p}^{2}/B_{\varphi}^{2}=\gamma_{\varphi\varphi}(\Omega_{\rm F}-\Omega)^{2}/\beta^{2}$, where $D_{\rm p}^{2}=D^{r}D_{r}+D^{\theta}D_{\theta}$ and $B_{\varphi}^{2}=B^{\varphi}B_{\varphi}$, as
$x=\left(1+C\frac{\Delta}{\Sigma}\right)\left(\frac{\Omega-\Omega_{\rm F}}{\Omega_{\rm H}-\Omega_{\rm F}}\right)^{2}.$ (25)
@@
This form of $x$ guarantees $\mathbf{D}=0$ where $\Omega=\Omega_{\rm F}$ (see Eqs. 9 and 16) and the Znajek condition (Eq. 14) at the horizon, where $\Delta=0$ and $\Omega=\Omega_{\rm H}$. We take the numerical constant $C=-1$. Even for the other values of $C$, we find that the solution hardly changes, since $\Delta/\Sigma\ll 1$ in the inflow region. This form of $x$ was found to be consistent with that seen in a numerical solution that takes account of the transverse force balance ogihara21 .
Figure 1: Solution of steady, axisymmetric MHD inflow towards the event horizon of a Kerr BH along a magnetic field line. The poloidal magnetic field is assumed to be in the split-monopole shape, i.e., $B^{\theta}=v^{\theta}=0$. The parameters are $a=0.8$, $\theta=\pi/3$, $\Omega_{\rm F}=0.5\Omega_{\rm H}$, $\mathcal{E}=-10$, and $\mathcal{L}\Omega_{\rm F}=-10.55$. The top left panel shows $M^{2}=4\pi\rho\alpha^{2}u_{\rm p}^{2}/B_{\rm p}^{2}$ with $-f$ and $-f+\beta^{2}$ as functions of $r-r_{\rm H}$, while the top right panel $-H_{\varphi}/4\pi|\eta|$, $-u_{t}$, and $u_{\varphi}$. The bottom left panel shows $-v^{r}$ and $v^{\varphi}$ with $\alpha/\sqrt{\gamma_{\varphi\varphi}}$, $\Omega$, and $\Omega_{\rm F}$ as functions of $r-r_{\rm H}$, while the bottom right panel $B_{\hat{r}}/4\pi|\eta|$, $-B_{\hat{\varphi}}/4\pi|\eta|$, $D_{\hat{\theta}}/4\pi|\eta|$, and $e/20\pi\eta^{2}$.
We set the parameter values as consistent with the GRMHD simulation results introduced in Section 1. The GRMHD simulations of RIAFs in the KS coordinates for $10<a<$ 1 with many different codes show similar quasi-steady structure of BH magnetospheres, which appears to be compatible with the current VLBI observational results nakamura18 ; porth19 ; narayan22 ; eht19b ; eht21 . In those numerical solutions, $\Omega_{\rm F}/\Omega_{\rm H}\sim 0.3-0.5$, and the poloidal magnetic field looks like a split-monopole type particularly at $r\lesssim 5$ where the RIAF can be geometrically thin due to pressure of the magnetosphere tchekhovskoy10 ; penna13 ; ripperda22 . The parameter ranges of $\mathcal{E}$ and $\mathcal{L}$ will be explained later.
@@
In Figure 1 we show an inflow solution for the parameters $\Omega_{\rm F}=0.5\Omega_{\rm H}$, $\mathcal{E}=-10$, and $\mathcal{L}\Omega_{\rm F}=-10.55$. We have assumed the poloidal magnetic field as a split-monopole type, $B^{\theta}=0$ (and thus $v^{\theta}=0$ from Eq. 18), and have fixed $\theta=\pi/3$ for solving the Bernoulli equation. The BH spin parameter is set to be $a=0.8$, for which the horizon radius and BH angular velocity are $r_{\rm H}\approx 1.57$ and $\Omega_{\rm H}\approx 0.255$, respectively. The upper left panel of Figure 1 shows Alfvén Mach number $M^{2}=4\pi\rho\alpha^{2}u_{\rm p}^{2}/B_{\rm p}^{2}$ as a function of $r-r_{\rm H}$. The inflow starts at the separation point, $r\simeq r_{\rm H}+1.3$, passes the inner Alfvén point where $M^{2}=-f$ and the inner magnetosonic point where $M^{2}=-f+\beta^{2}$, and continues towards the horizon.